OpenStudy (anonymous):

Please help me solve? I'm not getting it right... x^2+2x+9=0

OpenStudy (nubeer):

\[x = \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] This is the quadratic formula... from your question x^2+2x+9=0 a=1, b = 2,c = 9.. now plug in values u will get 2 values of x

OpenStudy (lacypennelll):